OpenStudy (anonymous):

how many zero of the function f(x)z^3+2)/z? and (2z-1)^7/z^3

OpenStudy (shamil98):

f(x) = z^3 + 2/z ?

OpenStudy (shamil98):

did you mean f(z)?...

OpenStudy (anonymous):

yes (z^3+z)/z

OpenStudy (shamil98):

set f(z) = 0 and solve for z.

OpenStudy (shamil98):

0 = z(z^2+1)/z 0 = z^2 + 1 ...

OpenStudy (shamil98):

and the same for the other one.

OpenStudy (anonymous):

i think we have two zero is that correct?

OpenStudy (shamil98):

yeah.

OpenStudy (anonymous):

but the answer in my book said no zero

OpenStudy (shamil98):

well its complex roots.

OpenStudy (shamil98):

z^2 + 1 = 0 z = +/- i

OpenStudy (anonymous):

yes but i confuse we the answer no zero